OpenStudy (anonymous):

Translate the word phrase into a math expression. 12 meters longer than his throw A. t ÷ 12 B. 12 · t C. t – 12 D. t + 12

OpenStudy (anonymous):

Help plzz

OpenStudy (anonymous):

D, let's say x is your answer, then the distance x = distance t + 12meters (that he threw more) so x = t+12. :) Hope I helped. Any questions?

OpenStudy (anonymous):

ty rws

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

Can you help me whith another one

OpenStudy (anonymous):

I'll try! :)

OpenStudy (anonymous):

Translate the word phrase into a math expression. Antwan pays $12 per hour to rent a canoe for the afternoon. A. 12 + h B. 12 · h C. 12 ÷ h D. h ÷ 12

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

B. Antwan will be paying $12 per hour and h equals hours. So 12·h = his total. Understand?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Good :) If you need any more help, I'm here!

OpenStudy (anonymous):

ok this one is tricky

OpenStudy (anonymous):

Which is an everyday situation for the math expression with the context provided? ; sharing something equally A. The Perez family drove 500 miles each day of their vacation. If their vacation lasted for n days, then is the total number of miles they traveled on their vacation. B. The community theater had a capacity of 500 seats. If n is the number of tickets sold for a performance, then is number of seats that were left empty for that performance. C. Jody had 500 dollars in her savings account. She withdrew n dollars to buy some camping supplies. She then had dollars left in her savings account. D. A store had 500 apples to be placed in bags with an equal number of apples in each bag. If n is the number of bags to be used, then is the number of apples in each bag.

OpenStudy (anonymous):

I'll try my best!

OpenStudy (anonymous):

Well.... A. = 500*n B. = 500-n C. = 500-n D. = 500/n It's D because it equally distributes the apples into each bag. Any more? :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem! :) That's what this site is for, right? If you need anymore help, just send me a message!

OpenStudy (anonymous):

Which is an everyday situation for the math expression with the context provided? 14 – g; comparing people's ages A. Jack is 14. Maura is g years younger than Jack. Jack's age is represented by 14 – g. B. Jack is 14. Maura is g years younger than Jack. Maura's age is represented by 14 – g. C. Jack is 14. Maura is g years older than Jack. Maura's age is represented by 14 – g. D. Jack's age is g. Maura is 14 years younger than Jack. Maura's age is represented by 14 – g.

OpenStudy (anonymous):

This is tricky for me lol

OpenStudy (anonymous):

Okay, this one is a bit trickier. It's defiantly not C or D. I'm kind of torn between a and b. I am going to go with A, as we know Jack is 14 and Maura is g. And he is represented.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which expression describes the total distance hiked for a person who hiked 6 miles a day for d days? A. 6d B. 6 + d C. D. d – 6

OpenStudy (anonymous):

What's C's suggestion?

OpenStudy (anonymous):



OpenStudy (anonymous):

sry

OpenStudy (anonymous):

6-d

OpenStudy (anonymous):

No need to be sorry. Okay, so this one is VERY easy. If someone walked 6 miles for 6 days and you wanted to find out how many miles they walked, the equation would be 6·d or 6d. :)

OpenStudy (anonymous):

6d?

OpenStudy (anonymous):

am i right?